ON THE A-X SYSTEM OF THE CALCIUM MONOMETHOXIDE RADICAL

P. CROZET, \underline{A. J. ROSS}, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France; C. LINTON, M. J. DICK, Physics Department, University of New Brunswick, Fredericton, NB, Canada E3B 5A3; A. G. ADAM and W. S. HOPKINS, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

Laser excitation spectra of the A ^{2}E ← X ^{2}A_{1} system of four isotopomers of calcium methoxide were recorded at the University of New Brunswick. The radicals were produced in a laser ablation source, using a 1.2 % mixture of methanol (^{12}CH_{3}OH, ^{13}CH_{3}OH, ^{12}CD_{3}OD and ^{13}CD_{3}OD) in He as a precursor. High resolution spectra were recorded in the range 15870 - 15985 cm^{−1} and 16320 - 16500 cm^{−1}(F.W.H.M. 250 MHz) using a tunable, single-mode cw dye laser (CR 699 + autoscan), operating with DCM and Kiton red dyes. The transitions were assigned to the two spin-orbit components A ^{2}E_{3/2} ← X ^{2}A_{1} and A ^{2}E_{1/2} ← X ^{2}A_{1} of the origin band, and of a band with v’_{4}(Ca-[OCH_{3}] stretch) = 1. Resolved J and K structures are observed. The analysis used an effective Hamiltonian given in the literature^{a}. The fits build on earlier work on the main isotopomer^{b}, and include spectroscopic data for the B ← X system given in the literature.^{c}^{d} We obtain structural parameters for the calcium monomethoxide radical from the rotational constants obtained from the different isotopomers.

aX. Liu, C.P. Damo, T.-Y. D. Liu, S. C. Foster, P. Misra, L. Yu and T. A. Miller J. Phys. Chem \underline{93} 2266-2275 (1989)

bP. Crozet, F. Martin, A.J. Ross, C. Linton, M.J. Dick and A.G. Adam J. Mol. Spectrosc. \underline{213} 28-34 (2002)

cC. J. Whitham, S. A. Beaton, Y. Ito and J. M. Brown J. Mol. Spectrosc. \underline{191} 286-294 (1998)

dK. C. Namiki, J. S. Robinson and T. C. Steimle J. Chem. Phys. \underline{13} 5283-5289 (1998)